OpenStudy (anonymous):

Let f(x)= x^2+6 and g(x)=x+8/x Find (g*f)(4)

OpenStudy (anonymous):

g(f(4))=? to find it,first find f(4) which is f(4)=4^2+6 f(4)=16+6 f(4)=22 then g(f(4))=g(22)=22+8/22 =22+0.36 =22.36

OpenStudy (anonymous):

ya got it?

OpenStudy (anonymous):

Or in fractions: \[\huge \color{green}{g(f(4)) = \frac{246}{11}}\]

OpenStudy (anonymous):

No ! lol. The answer choices are fractions so it makes it kind of difficult. Are you saying the anwser is 22.36 ?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

the fraction is 246/11

OpenStudy (anonymous):

But its not one of the anwser choices. you have 11/15 , 15/11 , 63/55 and 55/63

OpenStudy (anonymous):

as the answer choices

OpenStudy (anonymous):

Is it? \[(g \times f)(4)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

There is something wrong in the question I think..